OpenStudy (anonymous):

integral(1 to 4) (sqrt(t)ln(t))dx

OpenStudy (anonymous):

integration by parts :)

OpenStudy (unklerhaukus):

\[\int\limits_1^4\sqrt t\ln t\text dx\] \[=\sqrt t\ln t\int\limits_1^4\text dx\]

OpenStudy (unklerhaukus):

or did you mean dt

OpenStudy (anonymous):

yes dt sorry

OpenStudy (anonymous):

\[\int_{1}^{4} \sqrt{t} \ \ln{t} \ \text{d}t\]

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

well do integration by parts letting\[\text{d}v=\sqrt{t} \ \text{d}t\]\[u=\ln t\]

OpenStudy (anonymous):

du?

OpenStudy (anonymous):

\[\text{d}u=\frac{1}{t} \text{d}t \]???

OpenStudy (anonymous):

do u know integration by parts?

OpenStudy (anonymous):

um.. i got \[\frac{ 4 }{ 9 }(4\ln (64)-7)\]

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

let me check it

OpenStudy (anonymous):

thats right :)